# Can xLLMs Understand the Structure of Dialog? Exploring Multilingual Response Generation in Complex Scenarios

### v1

@@ Abstract
Multilingual research has garnered increasing attention, especially in the domain of dialogue systems. The rapid advancements in large language models (LLMs) have fueled the demand for high-performing multilingual models. However, two major challenges persist: the scarcity of high-quality multilingual datasets and the limited complexity of existing datasets in capturing realistic dialogue scenarios. To address these gaps, we introduce XMP, a high-quality parallel Multilingual(X) dataset sourced from Multi-party Podcast dialogues [^1] https://github.com/ahxchzt/XMP_data. Each sample in the dataset features at least three participants discussing a wide range of topics, including society, culture, politics, and entertainment. Through extensive experiments, we uncover significant limitations in previously recognized multilingual capabilities of LLMs when applied to such complex dialogue scenarios. For instance, the widely accepted multilingual complementary ability of LLMs is notably impacted. By conducting further experiments, we explore the mechanisms of LLMs in multilingual environments from multiple perspectives, shedding new light on their performance in real-world, diverse conversational contexts.
Can xLLMs Understand the Structure of Dialog? Exploring Multilingual Response Generation in Complex Scenarios
@@ 1 Introduction
Multilingual dialogue research has garnered significant attention Xue et al. (2021); Liu et al. (2020); Conneau et al. (2020), especially with the explosive growth of large language models (LLMs) like Touvron et al. (2023); Yang et al. (2024). The development of highly capable multilingual models is urgently needed, which undoubtedly requires high-quality multilingual dialogue datasets. However, while multilingual datasets have been developed for various NLP tasks Lin et al. (2021); Liu et al. (2023); Ponti et al. (2020); Muennighoff et al. (2022), there remains a stark scarcity of such datasets for dialogue generation tasks. The existing dialogue datasets are primarily focused on simple dialogue scenarios and are limited to a few high-resource languages, lacking support for complex multi-party, multi-turn dialogues Hu et al. (2019) and low-resource languages, as shown in Table 1. As a result, there exists a significant gap in the "multilingual + complex dialogue scenarios" domain, which has seriously hindered the development of multilingual dialogue. Furthermore, this gap raises doubts about the adaptability of current LLMs to multilingual settings in complex dialogue scenarios.
Our previous experiments revealed that current LLMs exhibit significant shortcomings in adapting to complex multilingual dialogue scenarios. Specifically, similar to the setting of Choenni et al. (2023), we conducted in-context learning (ICL) experiments using templates and examples in various languages for multi-party dialogue scenarios. We observed that, under various template configurations, increasing the number of examples often led to degraded performance—this was even true for English corpora. The phenomenon indicates that the large models themselves struggle with adapting to MCDS.
Dataset Task Type Sample Count Parallel Corpus Multilingual
UDC RG 320K No No
XCOPA MCQA 7.5k No Yes
XWinograd Reasoning 4.5K No Yes
XDailyDialog RR 52k Yes Yes
Table 1: Common multilingual or multi-party datasets. UDC refers to the Ubuntu Dialogue Corpus. Task types include RG for Response Generation, RR for Response Retrieval, and MCQA for Multiple-Choice Question Answering.
@@
To address this gap, we present XMP, a high-quality parallel multilingual dataset for multi-party dialogue. The dataset is built from real-world multi-speaker interactions in podcasts, and the original English data has been carefully translated into six languages: English, Chinese, Japanese, German, French, and Italian. The dataset consists of 312k conversations spanning diverse topics such as society, culture, politics, and entertainment. To ensure translation quality, we implemented a rigorous quality control process, including systematic data filtering and cleaning to guarantee the diversity and semantic coherence of the dialogue content.
Through extensive experiments, we found that several previously recognized multilingual capabilities Doddapaneni et al. (2021); Wang et al. (2019); Whitehouse et al. (2023) of LLMs no longer hold true on this dataset. For instance, the commonly accepted multilingual complementary ability—where combining English data with another language during fine-tuning significantly boosts the model’s performance on that language—was notably absent in our experiments. Specially, whether working with languages from the same language family (e.g., English + French) or different language families (e.g., English + Chinese), the results on our dataset were inferior to those obtained by training on a single language. Our key contributions are summarized as follows:
@@
  • We construct and release XMP, the first high-quality parallel multilingual dataset for multi-party dialogue, covering six languages and over 312k conversations. The dataset offers diverse topics and supports research into complex multilingual dialogue generation and cross-lingual transfer.
@@
  • We conduct extensive evaluations on XMP, revealing significant limitations in the multilingual capabilities of current LLMs, particularly their inability to leverage multilingual complementary effects in complex dialogue settings.
  • We provide detailed analysis and insights into the mechanisms of LLMs in multilingual environments, highlighting key challenges and opportunities for improving their adaptability to real-world, multi-party dialogue scenarios.
@@ 2.1 Multilingual Datasets for Dialogue
The development of high-quality multilingual dialogue datasets has significantly contributed to the advancement of multilingual research. In recent years, multilingual datasets have expanded to cover a wide range of tasks, such as XCOPA Ponti et al. (2020), which focuses on multiple-choice question answering, and XWinograd Muennighoff et al. (2022), which emphasizes reasoning tasks. However, dialogue-related datasets remain scarce and are often constrained to simple dialogue scenarios. For instance, XDailyDialog Liu et al. (2023) provides a multilingual parallel dialogue dataset with four languages, designed for response retrieval tasks. Similarly, XPersona Lin et al. (2021) extends the Persona-Chat dataset to six languages, focusing on cross-lingual personalized dialogue. While these datasets support tasks such as monolingual and cross-lingual dialogue, they are limited in complexity, often focusing on straightforward conversation patterns and neglecting multi-party and multi-turn dialogue structures. Moreover, they tend to overlook the crucial task of dialogue response generation, which is fundamental for building robust dialogue systems.
@@ 2.2 Multilingual Capabilities of Large Language Models
Recent advancements in large language models (LLMs), such as GPT-4 Achiam et al. (2023) and LLaMA Touvron et al. (2023), have showcased impressive multilingual capabilities. By leveraging instruction tuning, human feedback alignment Lai et al. (2024), and vast multilingual training corpora, these models have significantly improved performance across diverse languages, including many low-resource ones. However, their evaluation has predominantly focused on relatively structured tasks, rather than the more dynamic and context-rich settings of dialogue generation.
Multilingual complementarity, a key feature of cross-lingual transfer, has been extensively studied in the context of multilingual language models (MLMs). Choenni et al. (2023) investigated cross-lingual data sharing during fine-tuning and demonstrated that training samples from one language can reinforce or complement the knowledge acquired from other languages. Similarly, Doddapaneni et al. (2021) highlighted that multilingual models often achieve better performance by leveraging linguistic similarities across languages, enabling effective zero-shot and few-shot cross-lingual transfer. These findings suggest that MLMs can effectively exploit shared representations across languages for structured tasks, such as natural language inference and multiple-choice question answering.
However, despite these promising results, the adaptability of large language models (LLMs) to complex dialogue scenarios remains underexplored. Most existing research focuses on relatively structured tasks and overlooks dynamic settings like multi-turn, multi-party dialogue. Our findings indicate that multilingual complementarity often diminishes in these complex dialogue scenarios. For instance, training a model on both English and another language (e.g., French or Chinese) does not consistently outperform single-language training, even for linguistically similar languages.
@@ 3 Preliminary Experiments
To investigate the capabilities of large language models (LLMs) in generating responses for complex dialogue scenarios, we conducted preliminary experiments using approximately 1,000 English multi-party dialogue samples, each involving three or more speakers discussing topics such as culture, society, and entertainment. By testing various templates and different numbers of in-context learning (ICL) examples, we observed that LLMs failed to achieve significant improvements from ICL. Even for English, the language LLMs are typically most proficient in, the results showed almost no variation in performance as the number of ICL examples increased, with automatic metrics such as BLEU and ROUGE remaining nearly unchanged.
The phenomenon raised doubts about the ability of LLMs to handle complex dialogue scenarios. What is more, if their performance on English was already so limited, would it be even worse for other languages? To further explore this, we expanded the dataset by translating the dialogues into additional languages. Following the experimental setup in Choenni et al. (2023), we tested the multilingual data under diverse templates and varying ICL example counts. The results revealed a consistent performance decline for all non-English languages as the number of ICL examples increased, regardless of template design. The detailed experimental setup and results are provided in Appendix A.
These findings prompt deeper questions about the nature of LLMs’ capabilities and why ICL fails to yield improvements in this scenario. Much of LLMs’ strength comes from the immense volume of training data they are exposed to, enabling them to approximate conversational patterns and generate contextually plausible responses. However, in complex multi-party dialogue settings, where multiple speakers interact dynamically and linguistic nuances vary significantly across languages, the models’ reliance on data-driven pattern matching seems insufficient. They struggle to incorporate additional examples to refine their understanding of the dialogue’s structure or adapt their reasoning to the specific task context.
@@ 4 Data Collection
The Multilingual Complex Dialogue Dataset aims to provide high-quality parallel data for studying multilingual in complex multi-party scenarios. The raw data was collected from podcasts and subsequently translated into parallel corpora using Google Translate. In this section, we describe the process of collecting and preparing the dataset.
@@ 4.1 Data Source
In our data collection process, we chose podcasts as the primary data source due to their ability to provide high-quality transcripts and their frequent use of multi-party interaction formats involving three or more speakers. The format aligns well with the needs of multi-party dialogue research and offers a realistic representation of complex dialogue scenarios found in real-world interactions.
We collected data from various publicly available podcasts, including "Life Kit," "Pop Culture Happy Hour," "Short Wave," "Throughline," "Consider This from NPR," "It’s Been a Minute," "The NPR Politics Podcast," "The Lawfare Podcast," and "Planet Money." All these programs are publicly accessible and available for free download. Details regarding the crawling, filtering, cleaning, and segmentation processes are provided in Appendix B.
@@ 4.2 Translation Method
After completing the initial collection of English transcripts, the next challenge was expanding them into multiple languages without compromising contextual integrity or the complexity of the dialogue scenarios. The quality of large-scale parallel corpora depends heavily on the accuracy and coherence of the translation at the word, sentence, and conversation levels. To address this, we designed a comprehensive Translation Method Evaluation process to strike an optimal balance between translation quality and efficiency.
Translation Method W-Level U-Level C-Level
Google Translate 92.6 86.9 84.5
Baidu Translate 88.2 73.1 74.7
DeepL Translate 88.7 70.8 73.4
iFLYTEK Translate 86.2 70.2 72.3
LLaMA 3.1 90.4 80.3 82.8
ChatGPT-4o 94.5 85.9 87.8
Human Translation 99.2 98.5 95.4
Table 2: Evaluation Results of Translation Methods
@@
We began by selecting 50 representative samples that covered a diverse range of topics (e.g., law, politics, daily life), dialogue turns (ranging from 7 to 15 turns), and multi-party podcast segments. This ensured that the evaluation samples were sufficiently diverse in complexity, providing reliable references for subsequent large-scale translation.
For these samples, we employed several translation methods, including Google Translate, DeepL, Baidu Translate, iFLYTEK Translate, LLaMA 3.1, ChatGPT-4o and human translation. Each method translated the English transcripts into five other languages (Chinese, Japanese, French, German, and Italian), resulting in 300 translations (50 samples × 6 languages). To ensure objectivity and reliability, three linguistics experts reviewed each translation in multiple rounds. The review process identified issues such as lexical misusage, semantic distortion, or contextual inconsistency in the translated outputs. Ultimately, the evaluation focused on three dimensions:
  • Conversation-leve evaluation ensured contextual consistency and alignment with the original dataset.
As shown in Table 2, human translation consistently achieved the highest overall quality. However, due to the prohibitive cost and extended timeline of large-scale manual translation, we explored automated tools. Among Google Translate, DeepL, Baidu, iFLYTEK, LLaMA 3, ChatGPt-4o, both ChatGPT-4 and Google Translate have shown impressive results, with ChatGPT-4 slightly outperforming Google Translate. However, ChatGPT-4 requires careful handling of templates, as it may provide natural translations for some examples but sometimes offers literal translations that overlook specific contexts. Considering translation speed and cost-effectiveness, we opted for Google Translate as the more practical choice for large-scale multilingual tasks. Specific case comparisons of the translation tools are detailed in Appendix C.
4.3 Dataset Quality Analysis and Statistics
The dataset includes six languages: English, Chinese, Japanese, French, German, and Italian. While English serves as the original source, all other languages were generated using translation API. To systematically evaluate the quality of the dataset, we conducted a thorough assessment focusing on four key aspects: word-level accuracy, sentence fluency, semantic consistency, and contextual coherence.
To assess the quality of the dataset, we employed expert evaluation. Specially, we randomly selected 500 conversations, covering various topics. For each language, two evaluation experts were selected based on their proficiency in both English and the target language. The experts were trained on the evaluation process, including specific scoring criteria and illustrative examples for each rating level, to ensure consistency and objectivity. In cases where there was significant disagreement between the two experts’ scores, a third expert was consulted for adjudication. The final scores were aggregated and averaged to provide an overall quality score for each language. Detailed evaluation criteria and the scoring results for each language are provided in Appendix D.
@@ 4.3 Dataset Quality Analysis and Statistics
In Table 3, we present the detailed statistical analysis of the dataset. The dataset contains a total of 312k dialogues and 2,649k utterances, with an average of 8.47 turns per dialogue, indicating a high level of dialogue density. This ensures that each dialogue typically involves multiple exchanges, creating scenarios rich in multi-turn interactions. The dataset also emphasizes its multi-party characteristics, with an average of 3.64 speakers per dialogue. Notably, over 75% of the dialogues involve at least three speakers, as shown in Figure 1, reflecting the interaction between multiple distinct roles. This multi-party nature enhances the realism of the dataset and provides a robust foundation for studying complex dialogue scenarios. Moreover, the average number of turns per speaker is 2.33, indicating that most speakers contribute multiple times to the conversation. Additionally, we ensured that the target speaker (i.e., the speaker for whom a response is to be generated) had spoken at least once earlier in the dialogue history. This design guarantees that the model has sufficient contextual information to generate coherent and relevant responses.

### v2

# Can MLLMs Generalize to Multi-Party dialog? Exploring Multilingual Response Generation in Complex Scenarios

@@ Abstract
Current multilingual large language models(MLLMs) still focus on simple question-answering formats, often overlooking more complex dialogue scenarios. In other words, their capabilities of multilingual large models have yet to be validated in dialogue tasks with intricate structures. We therefore ask, Q1: How well do LLMs generalize to more complex dialog scenarios? Q2: Can supervised fine-tuning on a high-quality parallel benchmark restore this ability? Q3: Does the "multilingual complementarity" effect survive in the setting? To answer these questions, we introduce XMP [^1] https://anonymous.4open.science/r/XMP_data-5CC7, a high-quality parallel Multilingual(X) dataset sourced from Multi-party Podcast dialogues, which is the first parallel dataset focusing on multi-party dialogue scenarios. Most samples in the dataset feature three or more participants, discussing a wide range of topics. Through extensive experiments, we find that, R1: MLLMs fail to generalize to multi-party setting, R2: Fine-tuning on XMP improves only marginally, with the 70B model achieving at most a 1% absolute gain over its 8B counterpart; R3: Mixing languages during SFT is usually detrimental, with any benefits being marginal and limited to isolated cases in the 70B model.
Can MLLMs Generalize to Multi-Party dialog? Exploring Multilingual Response Generation in Complex Scenarios
@@ 1 Introduction
Amid the rising demand for seamless cross-lingual communication and the growing complexity of global language needs, multilingual research has garnered widespread attention Xue et al. (2021); Liu et al. (2020); Conneau et al. (2020a). Expectations for large language models (LLMs) have grown substantially, Touvron et al. (2023); Yang et al. (2024) particularly their ability to handle a wide range of multilingual NLP tasks Lin et al. (2021); Liu et al. (2023); Ponti et al. (2020); Muennighoff et al. (2022). Among these tasks, dialogue generation stands out as one of the most crucial. However, current LLMs have only demonstrated their multilingual capabilities in relatively simple, two-party dialogue scenarios. More complex interaction settings, such as multinational meetings or multi-player role-playing games are largely unexplored.
Figure 1: An example of a multi-party dialogue illustrating complex conversational dependencies, showcasing turn-taking, contextual references, and opinion shifts.
To date, no multilingual dataset exists that focuses specifically on more complex dialogue settings involving long contexts and multiple speakers Hu et al. (2019). An example of such a scenario is illustrated in Figure 1, where the structure and dynamics of the conversation highlight its complexity. This absence of suitable datasets creates a critical research gap, severely limiting our understanding of how effectively existing MLLMs generalize to realistic and complex dialogue scenario.
We systematically address the following clearly defined research questions, Q1: How well do LLMs generalize to more complex dialog scenarios? Q2: Can supervised fine-tuning on a high-quality parallel benchmark restore this ability? Q3: Prior studies like Mulcaire et al. (2019); Conneau et al. (2020b); Whitehouse et al. (2023) have shown that mixing multiple languages during training can enhance model performance. Does this effect still hold under the more complex dialogue settings considered in our work?
@@
To investigate these issues, we present XMP, a high-quality parallel multilingual dataset for multi-party dialogue. The dataset is built from real-world multi-speaker interactions in podcasts, and the original English data has been carefully translated into five languages: Chinese, Japanese, German, French, and Italian. The dataset consists of 312k conversations spanning diverse topics such as society, culture, politics, and entertainment. To ensure translation quality, we implemented a rigorous quality control process, including systematic data filtering and cleaning to maintain the diversity and semantic coherence of the dialogue content. In addition, we conducted a comprehensive quality assessment to ensure both cross-lingual consistency and cultural fidelity across all translated dialogues.
Through extensive experiments, we found that several previously recognized multilingual capabilities Doddapaneni et al. (2021); Wang et al. (2019); Whitehouse et al. (2023) of LLMs no longer hold true on this dataset. Specifically, R1: MLLMs exhibit significant limitations in generalizing to complex dialogue scenarios; R2: Fine-tuning on XMP improves only marginally, with the 70B model achieving at most a 1% absolute gain over its 8B counterpart; R3: The multilingual complementarity effect largely diminishes in complex, multi-party dialogues. Our key contributions are summarized as follows:
@@
  • We construct and release XMP, the first high-quality parallel multilingual dataset for multi-party dialogue, covering six languages and over 312k conversations. The dataset offers diverse topics and supports research into multilingual multi-party conversation and cross-lingual transfer.
@@
  • We conduct extensive evaluations on XMP, revealing significant limitations in the multilingual capabilities of current LLMs.
  • We reveal key insights into the performance of MLLMs, providing valuable guidance for advancing multilingual dialogue systems in realistic and challenging conversational contexts.
@@ 2.1 Multilingual Datasets for Dialogue
High-quality multilingual datasets play a crucial role in advancing multilingual NLP research, such as XCOPA Ponti et al. (2020) and XWinograd Muennighoff et al. (2022) primarily focus on structured tasks like multiple-choice question answering and reasoning. Although there exist multilingual dialogue datasets such as XDailyDialog Liu et al. (2023) and XPersona Lin et al. (2021), these datasets primarily emphasize straightforward two-party dialogues, typically lack complexity, rarely addressing multi-party scenarios, and frequently neglect the critical task of response generation. Given the fundamental importance of generating coherent and contextually relevant dialogue responses, this omission significantly restricts progress in multilingual dialogue system development.
2.2 Dialogue in Complex Scenarios
While dialogue generation has been extensively studied, most existing benchmarks focus on dyadic (two-party) conversations, which significantly limits their applicability to more realistic, multi-party scenarios.
Friends Shmueli and Ku (2019) is one of the most commonly used multi-party dialogue datasets, but it suffers from a key limitation: it includes only six fixed speakers from scripted television show dialogues. This restricts its generalizability, as it lacks the diversity and spontaneity found in real-world multi-party interactions.
Ubuntu Dialogue Corpus (UDC) Lowe et al. (2015) is another widely used dataset featuring technical support conversations from the Ubuntu community. However, the dialogues in UDC are primarily noisy forum threads with fragmented turn-taking and frequent context gaps. Moreover, the conversations are largely constrained to a narrow domain of technical troubleshooting, offering limited topic diversity.
@@ 2.3 Multilingual Capabilities of Large Language Models
Recent multilingual large language models (MLLMs), such as GPT-4 Achiam et al. (2023) and LLaMA Touvron et al. (2023), have shown strong multilingual capabilities Lai et al. (2024).
Further more, Cross-lingual data sharing has been extensively studied in the context of multilingual language models (MLMs). Choenni et al. (2023) investigated how training samples from one language can influence the learning of other languages during fine-tuning. Similarly, Doddapaneni et al. (2021) highlighted that multilingual models often achieve better performance by leveraging linguistic similarities across languages, enabling effective zero-shot and few-shot cross-lingual transfer.
Building upon these insights, we introduce the term Multilingual Complementarity to describe the phenomenon where multilingual models leverage data from multiple languages to enhance performance on a specific language task. While prior studies have observed this effect in structured tasks, its presence in complex generation tasks remains underexplored.
@@ 3 Preliminary Experiments
To investigate the capabilities of large language models (LLMs) in generating responses for complex dialogue scenarios, we conducted preliminary experiments using approximately 1,000 English multi-party dialogue samples. The detailed experimental setup and results are provided in Appendix A. By testing various templates and different numbers of in-context learning (ICL) examples, we observed that LLMs failed to achieve significant improvements from ICL. Even for English, the language LLMs are typically most proficient in, the results showed almost no variation in performance as the number of ICL examples increased.
The phenomenon raised doubts about the ability of LLMs to handle complex dialogue scenarios. What is more, if their performance on English was already so limited, would it be even worse for other languages? To further explore this, we expanded the dataset by translating the dialogues into additional languages. Following the experimental setup in Choenni et al. (2023), we tested the multilingual data under diverse templates and varying ICL example counts. The results revealed a consistent performance decline for all non-English languages as the number of ICL examples increased, regardless of template design.
These findings prompt deeper questions about the nature of LLMs’ capabilities and why ICL fails to yield improvements in this scenario. Much of LLMs’ strength comes from the immense volume of training data they are exposed to, enabling them to approximate conversational patterns and generate contextually plausible responses. However, in complex multi-party dialogue settings, where multiple speakers interact dynamically and linguistic nuances vary significantly across languages, the models’ reliance on data-driven pattern matching seems insufficient. They struggle to incorporate additional examples to refine their understanding of the dialogue’s structure or adapt their reasoning to the specific task context (R1).
@@ 4 Data Collection
To further evaluate the capabilities of MLLMs in multi-party dialogue and to support future related research, we propose XMP dataset.
@@ 4.1 Data Source
We collected data from various publicly available podcasts, including "Life Kit," "Pop Culture Happy Hour," "Short Wave," "Throughline," "Consider This from NPR," "It’s Been a Minute," "The NPR Politics Podcast," "The Lawfare Podcast," and "Planet Money." All these programs are publicly accessible and available for free download. Details regarding the crawling, filtering, cleaning, and segmentation processes are provided in Appendix B. We chose podcasts as the primary data source due to their ability to provide high-quality transcripts and their frequent use of multi-party interaction formats involving three or more speakers.
@@ 4.2 Translation Method
Singh et al. (2024) report that traditional machine translation systems often outperform large language models in translation quality, however, they tend to be less sensitive to cultural nuances. To validate which translation strategy is more suitable for our task, we evaluated multiple translation methods. Specifically, we experimented with commonly used tools in prior research Liu et al. (2023); Lin et al. (2021), including Google Translate, DeepL, Baidu Translate, and iFLYTEK Translate, alongside LLaMA 3.1 and ChatGPT-4o.
We adopted both human and automatic evaluations to ensure translation quality. For human evaluation, we selected 50 English dialogue samples spanning diverse topics and lengths (7–15 turns), and translated them into five languages (Chinese, Japanese, French, German, Italian). Three linguists reviewed each sample for word-level accuracy, fluency, and dialogue-level coherence. ChatGPT-4o and Google Translate received the highest overall ratings, though ChatGPT-4o occasionally introduced off-topic phrases (e.g., “Here is the translation…”), making it less suitable for large-scale batch translation. Given its consistency,cost, speed, and prior use in similar work, we selected Google Translate.
To further support this decision, we translated 5,000 samples into Chinese and German using both GPT-4o and Google Translate, and trained LLaMA 3.1 on each dataset. The generation outcomes showed minimal differences, with Google Translate slightly outperforming in some metrics. These results reinforce its suitability for scalable multilingual dialogue construction.
Full evaluation details and translation case analyses are provided in Appendix C
4.3 Dataset Quality Analysis
As mentioned before, Singh et al. (2024) purpose that machine translation systems tend to be less sensitive to cultural nuances. To systematically evaluate the quality of the whole dataset, we conducted a thorough assessment focusing on two key aspects: Cross-lingual Consistency Analysis and Linguistic and Cultural Consistency.
Dataset Task Type Multilingual Complex Context
UDC RG No Yes
XCOPA MCQA Yes No
XWinograd Reasoning Yes -
XDailyDialog RR Yes No
XMP(Ours) RG Yes Yes
Table 1: Common multilingual or multi-party datasets. Task types include RG for Response Generation, RR for Response Retrieval, and MCQA for Multiple-Choice Question Answering.
@@ 4.3 Dataset Quality Analysis
Specially, we randomly selected 1000 conversations (200 samples for each language), covering various topics. For each language, two evaluation experts were selected based on their proficiency in both English and the target language. The experts were trained on the evaluation process, including specific scoring criteria and illustrative examples for each rating level, to ensure consistency and objectivity. In cases where there was significant disagreement between the two experts’ scores, a third expert was consulted for adjudication. The final scores were aggregated and averaged to provide an overall quality score for each language. Detailed evaluation criteria and the scoring results for each language are provided in Appendix D. Overall, we found that the dataset demonstrated good quality across the board, achieving strong results in both Cross-lingual Consistency and Linguistic and Cultural Consistency, indicating that the translated texts can preserve semantic consistency while also taking cultural differences into account to some extent. Although relatively low-resource languages such as Italian received slightly lower scores than others, their performance still fell within an acceptable range.
4.4 Dataset Statistics
In Table 1, we compare our XMP dataset with several commonly used multilingual or multi-party datasets. XMP is the first high-quality parallel multilingual dataset for multi-party dialogue.
@@ 4.4 Dataset Statistics
In Table 2, we present key statistics of the dataset, which includes 312k dialogues and 2,649k utterances, averaging 8.47 turns per dialogue. This high dialogue density supports rich multi-turn interactions. On average, each dialogue features 3.64 speakers, and over 75% involve at least three participants (Figure 2), highlighting the dataset’s multi-party nature and the diverse role dynamics it captures.
Each speaker contributes an average of 2.33 turns, suggesting active participation across the dialogue. Additionally, we ensured that the target speaker (i.e., the speaker for whom a response is to be generated) had spoken at least once earlier in the dialogue history. This design guarantees that the model has sufficient contextual information to generate coherent and relevant responses.